OpenStudy (anonymous):

Determine the intercepts and use symmetry to assist in sketching the graph of the equation y=3x-x^2

OpenStudy (anonymous):

rewrite: y= -x^2 +3x x symmetry = -3/(-2) = 3/2 y = -(3/2)^2 + 3(3/2) = (-9/4) + (9/2) = 9/4 vertex: ( 3/2, 9/4 ) if x= 0 , y intercept:=0 -x^2 +3x = 0 -x(x -3) x intercept : x= 3 or x= 0